A perpetuity of $5,000 per year beginning today offers a 15% return. What is its present value?

Multiple Choice

$33,333.33

$37,681.16

$65,217.39

$38,333.33
Answer #1

Present value of perpetuity paid at beginning is

= C(1+r)/r

Where c is periodic payment

R is rate

Pv = 5000(1.15)/15% =38333.33

Option 1 is correct